### v1

# InstructLayout: Instruction-Driven 2D and 3D Layout Synthesis with Semantic Graph Prior

Abstract
Comprehending natural language instructions is a charming property for both 2D and 3D layout synthesis systems. Existing methods implicitly model object joint distributions and express object relations, hindering generation’s controllability. We introduce InstructLayout, a novel generative framework that integrates a semantic graph prior and a layout decoder to improve controllability and fidelity for 2D and 3D layout synthesis. The proposed semantic graph prior learns layout appearances and object distributions simultaneously, demonstrating versatility across various downstream tasks in a zero-shot manner. To facilitate the benchmarking for text-driven 2D and 3D scene synthesis, we respectively curate two high-quality datasets of layout-instruction pairs from public Internet resources with large language and multimodal models. Extensive experimental results reveal that the proposed method outperforms existing state-of-the-art approaches by a large margin in both 2D and 3D layout synthesis tasks. Thorough ablation studies confirm the efficacy of crucial design components.
Index Terms:
layout synthesis, controllable generation, graph diffusion models, scene graphs.
1 Introduction
[…]
[lines=2]Automatically synthesizing controllable and appealing 2D and 3D layouts has been a challenging task for computer vision and graphics [1; 2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12; 13; 14; 15], which requires building a generative model by multimodal conditional information. An ideal layout synthesis system should fulfill at least three objectives: (1) comprehending instructions in natural languages, thus providing an intuitive and user-friendly interface; (2) designing object compositions that exhibit aesthetic appeal and thematic harmony; (3) placing objects in appropriate positions adhering to their functions and regular arrangements.
To enhance the controllability of layout synthesis, we need to address natural language instructions as inputs. However, natural instructions for layout design often rely on abstract object relationships, which has posed a significant challenge for recent advancements in 2D and 3D layout synthesis. Previous works [16; 17; 18; 19; 20; 21] primarily exclude instructional prompts or do not prioritize them, leading to a low level of controllability and interpretability in the generated layouts. Other studies [22; 23; 24] utilize relation graphs to provide explicit control over object interactions, which are, however, too complicated and fussy for human users to specify. Moreover, previous works merely represent objects by categories  [22; 6; 19; 17] or low-dimensional features [25; 11; 12; 21] which lack visual appearance details, resulting in style inconsistency and constraining customization options in layout synthesis.
To address these issues, we present InstructLayout, a novel generative framework for both 2D and 3D layout synthesis with natural language instructions. Figure 1 illustrates the overview of proposed method. InstructLayout basically comprises two parts: a semantic graph prior and a layout decoder. In the first stage, it takes instructions about partial object arrangement and attributes as conditions and learns the conditional distribution of semantic graphs for the holistic layout. In the second stage, harnessing the well-structured and informative graph latents, the layout decoder can easily generate 2D poster and 3D scene layouts that exhibit semantic consistency while closely adhering to the provided instructions. For 2D poster layout synthesis, we additionally provide a tagline generator to fill the bounding boxes with text content for real-world applications. With the learned semantic graph prior, InstructLayout also achieves a wide range of instruction-driven generative tasks in a zero-shot manner.
Figure 1: Method overview. For both 3D scene and 2D poster layout synthesis tasks, (1) InstructLayout first designs a holistic semantic graph based on user instructions. Within the graph, each node is an object endowed with semantic features such as categories and appearances, and each edge represents a spatial relationship between objects. (2) It proceeds to arrange objects in a scene or canvas by decoding spatial attributes from the informative graph prior.
Specific conditional diffusion models are devised for both parts of InstructLayout. Benefitting from the two-stage scheme, it can separately handle discrete and continuous attributes of layouts, drastically reducing the burden of network optimization. To enhance the capability of aesthetic design, InstructLayout utilizes the naturally discrete features of 2D objects, such as fonts, colors, and shapes. In contrast, in terms of 3D objects, it leverages object geometrics and appearances by quantizing semantic features from a multimodal-aligned model [26]. Consequently, our model can handle these discrete features in a unified manner.
To fit practical scenarios and promote the benchmarking of instruction-drive layout synthesis, we curate two high-quality datasets containing paired layouts and instructions with the help of large language and multimodal models [27; 28; 29] for both 2D E-commerce poster and 3D indoor scene synthesis tasks. Comprehensive quantitative evaluations reveal that InstructLayout surpasses previous state-of-the-art methods by a large margin in terms of both generation controllability and fidelity. Each essential component of our method is carefully verified through ablation studies.
Our contributions can be summarized as follows:
  • We present a unified instruction-driven generative framework for both 3D scene and 2D poster layout synthesis that integrates a semantic graph prior and a layout decoder to improve fidelity and controllability.
  • The proposed semantic graph prior jointly models object attributes and layout distributions, facilitating various downstream applications in a zero-shot manner.
  • We curate two high-quality datasets to promote the benchmarking of instruction-driven 2D and 3D layout synthesis, and quantitative experiments demonstrate that the proposed method significantly outperforms existing state-of-the-art techniques.
2 Related Work
2.1 3D Scene Layout Synthesis
Graphs have been used to guide complex scene synthesis in the form of scene hierarchies [30; 31], parse trees [32], scene graphs [33; 34], etc. Wang et al. [25] utilize an image-based module and condition its outputs on the edges of a relation graph within each non-differentiable step. They also adopt an autoregressive model [35] to generate relation graphs, which are unconditional and have limited object attributes. Other works [22; 23; 24] adopt conditional VAEs [36; 37] with graph convolutional networks [38] to generate layouts. While offering high controllability, these methods demand the specification of elaborate graph conditions, which are notably more intricate than those driven by natural languages. With the advent of attention mechanisms [39], recent approaches [16; 17; 20; 21] can implicitly acquire object relations by self-attention and condition scene synthesis with texts by cross-attention. However, text prompts in these works tend to be relatively simple, containing only object categories or lacking layout descriptions, limiting the expressiveness and customization. Implicit relation modeling also significantly hinders their controllability.
2.2 2D Poster Layout Synthesis
Numerous methods have been proposed to automatically synthesize 2D layouts for posters, user interfaces, E-commerce advertisements, etc. Prior research [7; 6; 40; 11; 19] has primarily concentrated on generating layouts from scratch on a blank canvas, while others [18; 41; 42; 10; 12] focus on image content-aware layout synthesis, leading to more practical applications. However, they are either confined to unconditional generation based solely on the canvas(blank or image prior) [7; 6; 41; 42] or relying on object categories [11; 19], lacking the capability to control the layout synthesis process using user-friendly instructions. Furthermore, due to the scarcity of high-quality datasets, only a few studies [12; 43] have delved into predicting 2D object attributes and their textual content, which is vital for generating realistic and informative layouts.
2.3 Generative Models for Graphs
There have been lots of endeavors on generative models for undirected graphs, molecules and scene graphs by autoregressive models [44; 45], VAEs [46; 47], GANs [48; 49] and diffusion models [50; 51; 52; 53]. However, none of them is capable of understanding text conditions. Longland et al. [54] and Lo et al. [55] employ VAE and GAN, respectively, for text-driven generation of simple undirected graphs without semantics. We present a pioneering effort to generate holistic semantic graphs with expressive instructions.
3 Method
3.1 Problem Statement
Denote $\mathcal{S}\coloneqq\{\mathcal{S}_{1},\dots,\mathcal{S}_{M}\}$ as a collection of 2D or 3D layouts. Each layout $\mathcal{S}_{i}$ is composed of multiple objects $\mathcal{O}_{i}\coloneqq\{\mathbf{o}_{j}^{i}\}_{j=1}^{N_{i}}$ with distinct attributes $\mathbf{o}_{j}^{i}\coloneqq\{c_{j}^{i},\mathbf{s}_{j}^{i},\mathbf{f}_{j}^{i}\}$, including category $c_{j}^{i}\in\{1,...,K_{c}\}$, where $K_{c}$ is the number of object classes in $\mathcal{S}$, spatial feature $\mathbf{s}_{j}^{i}$, such as scale and location, and semantic feature $\mathbf{f}_{j}^{i}$ like appearance and style. To set up a layout, one can generate each object (a 3D model or tagline text) or retrieve it from a database based on $c$ and $\mathbf{f}$. These objects are then resized and transformed to the expected coordinates by corresponding $\mathbf{s}$.
Given instructions $\mathbf{y}$, our goal is to learn the conditional scene distribution $q(\mathcal{S}|\mathbf{y})$. Rather than direct modeling [17; 19; 12; 21], we employ well-structured and informative graphs to serve as general and semantic latent. Each graph $\mathcal{G}_{i}$ contains a node set $\mathcal{V}_{i}\coloneqq\{\mathbf{v}_{j}^{i}\}_{j=1}^{N_{i}}$ and a directed edge set $\mathcal{E}_{i}\coloneqq\{e_{jk}^{i}|\mathbf{v}_{j}^{i},\mathbf{v}_{k}^{i}\in\mathcal{V}_{i}\}$. A node $\mathbf{v}_{j}^{i}$ functions as a high-level representation of an object $\mathbf{o}_{j}^{i}$, and a directed edge $e_{jk}^{i}$ explicitly conveys the relations between objects.
To this end, we propose a generative framework, InstructLayout, that consists of two components: (1) semantic graph prior $p_{\phi}(\mathcal{G}|\mathbf{y})$ (Sec. 3.3) that jointly models high-level object and relation distributions conditioned on $\mathbf{y}$; (2) layout decoder $p_{\theta}(\mathcal{S}|\mathcal{G})$ (Sec. 3.4) that produces precise layout configurations with semantic graphs prior. Since $\mathcal{G}$ is deterministic by corresponding $\mathcal{S}$, the two networks together yield an instruction-driven generative model for 2D and 3D layouts:
$p_{\phi,\theta}(\mathcal{S}|\mathbf{y})=p_{\phi,\theta}(\mathcal{S},\mathcal{G}|\mathbf{y})=p_{\phi}(\mathcal{G}|\mathbf{y})p_{\theta}(\mathcal{S}|\mathcal{G}).$ (1)
3.2 Preliminary: General Diffusion Models
Diffusion generative models [56] consist of a non-parametric forward process and a learnable reverse process. The forward process progressively corrupts a data point from $q(\mathbf{x}_{0})$ to a sequence of increasingly noisy latent variables: $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$. A neural network is trained to reverse the process by denoising them iteratively: $p_{\psi}(\mathbf{x}_{0:T}|\mathbf{c})=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\psi}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})$, where $\mathbf{c}$ is an optional condition to guide the reverse process as needed. These two processes are supposed to admit $p(\mathbf{x}_{T})\approx q(\mathbf{x}_{T}|\mathbf{x}_{0})$ for a sufficiently large $T$. The generative model is optimized by minimizing a variational upper bound on $\mathbb{E}_{q(\mathbf{x}_{0})}\left[-\log p_{\psi}(\mathbf{x}_{0})\right]$:
$\mathcal{L}_{\text{vb}}\coloneqq\mathbb{E}_{q(\mathbf{x}_{0})}[\ \ \underbrace{D_{\text{KL}}[q(\mathbf{x}_{T}|\mathbf{x}_{0})\|p(\mathbf{x}_{T})]}_{\mathcal{L}_{T}}\\ +\sum_{t=2}^{T}\mathcal{L}_{t-1}\underbrace{-\mathbb{E}_{q(\mathbf{x}_{1}|\mathbf{x}_{0})}[\log p_{\psi}(\mathbf{x}_{0}|\mathbf{x}_{1},\mathbf{c})]}_{\mathcal{L}_{0}}\ \ ],$ (2)
where $\mathcal{L}_{t-1}\coloneqq D_{\text{KL}}[q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})\|p_{\psi}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})]$ and $\mathcal{L}_{T}$ is constant during training so can be ignored. $D_{\text{KL}}[\cdot\|\cdot]$ indicates the KL divergence between two distributions.
Diffusion models can accommodate both continuous and categorical distributions by a meticulous design of the diffusion kernel $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$. A prominent choice for continuous variables is the Gaussian kernel [57; 58; 21]. Uniform transition is widely adopted for categorical variables [59; 60; 52]. In this work, we utilize both the discrete and continuous diffusion models for the two stages in InstructLayout respectively to model distinct layout attributes.
3.3 Semantic Graph Prior
A semantic graph $\mathcal{G}$ represents an abstract structure of the underlying layout by a node set and an edge set. Each node includes the category $c$ and semantic features $\mathbf{f}$ of an object. Spatial features $\mathbf{s}$ can be derived from relations represented by edges, so we leave them to the decoder $p_{\theta}(\mathcal{S}|\mathcal{G})$.
Figure 2: Semantic Graph Prior (Sec. 3.3). (a) 3D Feature Discretization (Sec. 3.3.1). Unlike the naturally discrete features of 2D poster graphics, semantic features for 3D objects are extracted from a frozen multimodal-aligned model OpenShape and then quantized by codebook entries. (b) Discrete Semantic Graph Diffusion (Sec. 3.3.2). Three categorical variables in $G_{0}$ are independently diffused; Empty states are not depicted for concision; A graph Transformer with a frozen text encoder and optional image encoder for 2D poster learns the semantic graph prior by iteratively denoising corrupted graphs.
3.3.1 Feature Discretization
For the 2D poster layout synthesis task, we consider graphic features, including types of text fonts, italicized or not, and border shapes, which are naturally discrete variables like category. For the colors of text, border, and gradation that are in continuous values, we discretize them in the CIE Lab colorspace following Zhang et al. [61] to enhance color saturation and constrain the feature space.
While for the 3D scene layout synthesis task, it’s hard to describe geometric shapes and visual appearance of 3D models by explicit categorical attributes, so we utilize a pretrained multimodal-aligned model OpenShape [26] to extract their semantic features. However, these high-dimensional features (e.g., $d=1280$ in OpenShape pointbert-vitg14) are too complicated to model. We circumvent this drawback by introducing a vector-quantized variational autoencoder (VQ-VAE) [62; 63] for feature vectors and quantize each of them by several learnable tokens from the maintained codebook, as depicted in Figure 2(a). The intuition behind it is that 3D models share general intrinsic characteristics, encompassing attributes like colors, materials, and basic geometric shapes. Indexing semantic features in a discrete space achieves a compact feature space for distribution modeling and dramatically reduces the cost of operating in a continuous space.
3.3.2 Discrete Semantic Graph Diffusion
After feature discretization, all attributes in a semantic graph are categorical: a graph $\mathcal{G}_{i}\coloneqq(\mathcal{C}_{i},\mathcal{F}_{i},\mathcal{E}_{i})$ with $N_{i}$ nodes, where $\mathcal{C}_{i}\coloneqq\{1,\dots,K_{c}\}^{N_{i}}$, $\mathcal{E}_{i}\coloneqq\{1,\dots,K_{e}\}^{N_{i}\times N_{i}}$ and $\mathcal{F}_{i}\coloneqq\{1,\dots,K_{f}\}^{N_{i}\times n_{f}}$, where $n_{f}$ is the number of discrete features. While it is possible to embed discrete variables in continuous spaces using one-hot encodings, it diminishes the sparsity inherent in the original data and imposes a substantial burden on network optimization. Instead, we propose to model the semantic graph prior by a mask-based discrete diffusion model.
For a scalar discrete random variable with $K$ categories $x\in\{1,\dots,K\}$, diffusion noise is defined by a series of transition matrices $\mathbf{Q}\in\mathbb{R}^{K\times K}$. The forward process at timestep $t$ is expressed as $q(x_{t}|x_{t-1})\coloneqq\mathbf{x}_{t}^{\top}\mathbf{Q}_{t}\mathbf{x}_{t-1}$, where $\mathbf{x}_{t}\in\mathbb{R}^{K}$ is the column one-hot encoding for $x_{t}$ and $[\mathbf{Q}_{t}]_{mn}\coloneqq q(x_{t}=m|x_{t-1}=n)$ is the probability that $x_{t-1}$ transits to the category $m$ from $n$. The probabilistic distribution of $x_{t}$ can be directly derived from $x_{0}$: $q(x_{t}|x_{0})\coloneqq\mathbf{x}_{t}^{\top}\bar{\mathbf{Q}}_{t}\mathbf{x}_{0}$, where $\bar{\mathbf{Q}}_{t}\coloneqq\mathbf{Q}_{t}\cdots\mathbf{Q}_{1}$.
Instead of commonly used Gaussian or uniform transitions for graph generation [50; 59; 51; 52], we propose to diffuse semantic graphs by independently masking graph attributes (i.e., object class $c$, discrete feature indices $f$ and relation $e$) by introducing an absorbing state [MASK] [60; 64] to each uniform transition matrix. For object class $c$, its transition matrix is defined as:
$\mathbf{Q}_{t}^{\mathbf{C}}\coloneqq\left[\begin{array}[]{ccccc}\alpha_{t}^{c}+\beta_{t}^{c}&\beta_{t}^{c}&\cdots&\beta_{t}^{c}&0\\ \beta_{t}^{c}&\alpha_{t}^{c}+\beta_{t}^{c}&\cdots&\beta_{t}^{c}&0\\ \vdots&\vdots&\ddots&\beta_{t}^{c}&0\\ \beta_{t}^{c}&\beta_{t}^{c}&\beta_{t}^{c}&\alpha_{t}^{c}+\beta_{t}^{c}&0\\ \gamma_{t}^{c}&\gamma_{t}^{c}&\gamma_{t}^{c}&\gamma_{t}^{c}&1\end{array}\right],$ (3)
by which $c_{t}$ has a probability of $\gamma_{t}^{c}$ to be masked, a probability of $\alpha_{t}^{c}$ to maintain the same, leaving a chance of $1-\gamma_{t}^{c}-\alpha_{t}^{c}$ for uniform sampling. [MASK] will always stay in its state. Transition matrices for $f$ and $e$, denoted as $\mathbf{Q}_{t}^{\mathbf{F}}$ and $\mathbf{Q}_{t}^{\mathbf{E}}$ respectively, exhibit analogous structures. Schedules of $(\alpha_{t},\beta_{t},\gamma_{t})$ are designed to admit that the initial states for semantic graphs are all masked.
Since the number of objects varies across different layouts, semantic graphs are padded by empty states to maintain a consistent number of $N$ objects. One-hot encodings for scalar variables $c$, $f$ and $e$ in a layout are denoted as $\mathbf{C}\in\mathbb{R}^{N\times(K_{c}+2)}$, $\mathbf{F}\in\mathbb{R}^{N\times n_{f}\times(K_{f}+2)}$ and $\mathbf{E}\in\mathbb{R}^{N\times N\times(K_{e}+2)}$ respectively. Here, “$+2$” accounts for each variable’s two extra states (i.e., empty state and mask state). A one-hot encoded semantic graph $G_{0}\coloneqq(\mathbf{C}_{0},\mathbf{F}_{0},\mathbf{E}_{0})$ at timestep $t$ is formulated as
$q(G_{t}|G_{0})=(\bar{\mathbf{Q}}_{t}^{\mathbf{C}}\mathbf{C}_{0},\bar{\mathbf{Q}}_{t}^{\mathbf{F}}\mathbf{F}_{0},\bar{\mathbf{Q}}_{t}^{\mathbf{E}}\mathbf{E}_{0}).$ (4)
The process for learning the graph prior is illustrated in Figure 2(b). The independent diffusion with mask states offers two significant advantages:
  • Perturbed states for one variable (e.g., $\mathbf{C}$) could be recovered by incorporating information from uncorrupted portions of the other variables (e.g., $\mathbf{F}$ and $\mathbf{E}$), compelling the semantic graph prior to learning from the interactions among different scene attributes.
  • The introduction of mask states facilitates the distinction between corrupted and clean states, thus simplifying the denoising task.
These benefits are critical, especially for intricate semantic graphs and diverse downstream generative tasks, compared with simple graph generative tasks [50; 51; 52]. An ablation study on the choice of $\mathbf{Q}$ is provided in Sec. 6.2.
The output of the graph prior network is re-parameterized to produce the clean scene graphs $\hat{G}_{0}$, which is then diffused to get the predicted posterior for computing the variational bound in Equation 2: $p_{\phi}(G_{t-1}|G_{t},\mathbf{y})\propto\sum_{\hat{G}_{0}}q(G_{t-1}|G_{t},\hat{G}_{0})p_{\phi}(\hat{G}_{0}|G_{t},\mathbf{y})$. Training objective for $p_{\phi}$ is a weighted summation of variational bounds for three random variables conditioned on $\mathbf{y}$:
$\mathcal{L}_{\text{vb}}^{\mathcal{G}|\mathbf{y}}\coloneqq\mathcal{L}_{\text{vb}}^{\mathbf{C}|\mathbf{y}}+\lambda_{f}\cdot\mathcal{L}_{\text{vb}}^{\mathbf{F}|\mathbf{y}}+\lambda_{e}\cdot\mathcal{L}_{\text{vb}}^{\mathbf{E}|\mathbf{y}},$ (5)
where $\lambda_{e}$ and $\lambda_{f}$ are hyperparameters to adjust the relative importance of three components in the semantic graph.
3.4 Layout Decoder
Figure 3: (a) Layout Decoder (Sec. 3.4). Gaussian noises are sampled to attach at every node of semantic graphs; A graph Transformer processes these graphs iteratively to remove noises and generate spatial features (Sec. 3.4.1); A tagline generator followed to fill textual contents in each bounding box for 2D poster layout synthesis (Sec. 3.4.2). (b) Graph Transformer (Sec. 3.5). A graph Transformer consists of a stack of $M$ blocks, each comprising graph attention, MLP, and optional cross-attention modules for conditions; AdaLN and multihead scheme are not depicted for concision.
3.4.1 Spatial Feature Generation
Modeling layouts’ spatial features become easy with the semantic graph prior. Spatial features basically include two attributes $\mathbf{s}\coloneqq\{\mathbf{l},\mathbf{m}\}$, where $\mathbf{l}\in\mathbb{R}^{2}$ or $\mathbb{R}^{3}$ is the location of an object in 2D or 3D space and $\mathbf{m}\in\mathbb{R}^{2}$ or $\mathbb{R}^{3}$ is its scale. In 3D scenes, an object has an orientation attribute $r\in\mathbb{R}$, allowing it to rotate along the gravity axis. We parameterize the $SO(2)$ rotation by $[\cos(r),\sin(r)]^{\top}$ to continuously represent $r$ [65]. Consequently, the spatial features of each layout $\mathcal{S}_{i}$ can be expressed as matrices $\mathbf{L}_{i}\in\mathbb{R}^{N_{i}\times d_{l}}$, where $d_{l}=4$ for 2D posters and $d_{l}=8$ for 3D scenes. Note that $\mathcal{S}=(\mathbf{L},\mathcal{G})$, so generating layouts $p_{\theta}(\mathcal{S}|\mathcal{G})$ is equivalent to learning the conditional distributions of layout configurations $p_{\theta}(\mathbf{L}|\mathcal{G})$. For the 2D layout synthesis task, although spatial relations can be inferred from edges in the semantic graph, we additionally provide clean product poster images $\mathbf{y}_{\text{image}}$ as conditions for more precise arrangements and to avoid occluding the principal products.
A continuous diffusion model with variance-preserving Gaussian kernels [57; 58] is adopted to learn $p_{\theta}(\mathbf{L}|\mathcal{G})$. Following Ho et al. [57], the variational bound in Equation 2 for the decoder $p_{\theta}(\mathbf{L}|\mathcal{G})$ is reweighted and simplified:
$\mathcal{L}_{\text{simple}}\coloneqq\mathbb{E}_{\mathbf{L}_{0},t,\boldsymbol{\epsilon}}\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{L}_{t},t,\mathcal{G}\right)\|^{2},$ (6)
where $t$ is sampled from a uniform distribution $\mathcal{U}(1,T)$, $\boldsymbol{\epsilon}$ is sampled from a standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\mathbf{L}_{t}$ is the noised spatial features. A diagram of the layout decoder is depicted in Figure 3(a). Intuitively, the network $\boldsymbol{\epsilon}_{\theta}$ is trained to predict noise $\boldsymbol{\epsilon}$ in $\mathbf{L}_{t}$.
3.4.2 Tagline Generation for 2D Poster
After synthesizing semantic features $\mathbf{f}$ and spatial features $\mathbf{s}$, the pipeline for 3D scene synthesis is completed by retrieving 3D objects from an off-the-shelf dataset based on the generated $\mathbf{f}$ and transforming them to appropriate positions using the generated $\mathbf{s}$.
For 2D poster synthesis, we proceed to predict the textual contents in bounding boxes using all generated $\mathbf{f}$ and $\mathbf{s}$ associated with user-provided product descriptions and images. Concretely, we primarily adopt the method proposed by Gao et al. [43] to generate tagline captions. This method involves an autoregressive transformer decoder considering multimodal conditions such as background images, product descriptions, and spatial features. A progressive training strategy trains it through two tasks, including caption generation and caption matching. However, the absence of clean product images and semantic features of text bounding boxes results in suboptimal tagline generation. To address this issue, we utilize high-quality cleaned product images processed according to the strategy proposed in Sec. 5.1 and incorporate the generated semantic features from Sec 3.3 as an additional robust condition to guide generation.
3.5 Model Architecture
The general-purpose Transformer [39] architecture is used for all models across tasks in this work.
Vanilla Transformer. As illustrated in Figure 2(a), $n_{f}$ learnable tokens are employed with a stack of cross-attentions to extract information from object semantic features $\mathbf{f}$ in the encoder $E$ in VQ-VAE. Regarding the decoder $D$, $n_{f}$ vectors retrieved from the codebook $\mathcal{Z}$ are fed to another Transformer, and an average pooling on the top of it is applied to aggregate information.
Graph Transformer. The semantic graph prior and spatial decoder share the same model architecture as shown in Figure 3(b). Since relation $e_{jk}$ can be determined by $e_{kj}$, only the upper triangular part of the relation matrix is necessary. Object categories and features together form input tokens for Transformers. Message passing on graphs is operated via node self-attention and node-edge fusion with FiLM [66], which linearly modulates edge embeddings and node attention matrices before softmax [67; 52]. Timestep for diffusion $t$ is injected by AdaLN [68; 69]. In the graph prior $p_{\phi}(\mathcal{G}|\mathbf{y})$, instructions and optional product images are embedded by a frozen multimodal encoder CLIP [70] and consistently influence network outputs by cross-attention mechanisms. Layout decoder $p_{\theta}(\mathcal{S}|\mathcal{G})$ is conditioned on semantic graphs by appending sampled Gaussian noises on node embeddings, which are then iteratively denoised to produce layout spatial features.
Permutation Non-invariance. Although $\mathcal{G}$ should ideally remain invariant to node permutations, invariant diffusion models could encounter learning challenges for multi-mode modeling. Thus, each node feature is added with positional encodings [39; 71; 21] before the permutation-equivariant Transformer. Exchangeability for graph prior distributions is strived by random permutation augmentation during the training process. Ablation on the permutation non-invariance is provided in Sec. 6.2.
4 InstructLayout for 3D Scene Synthesis
[…]
Bedroom ATISS 48.13$\pm$2.50 119.73$\pm$1.55 6.95$\pm$0.06 0.39$\pm$0.02 59.17$\pm$1.39
DiffuScene 56.43$\pm$2.07 123.09$\pm$0.79 7.13$\pm$0.16 0.39$\pm$0.01 60.49$\pm$2.96
Ours 73.64$\pm$1.37 114.78$\pm$1.19 6.65$\pm$0.18 0.32$\pm$0.03 56.02$\pm$1.43
Living room ATISS 29.50$\pm$3.67 117.67$\pm$2.32 6.08$\pm$0.13 17.60$\pm$2.65 69.38$\pm$3.38
DiffuScene 31.15$\pm$2.49 122.20$\pm$1.09 6.10$\pm$0.11 16.49$\pm$1.24 72.92$\pm$1.29
Ours 56.81$\pm$2.85 110.39$\pm$0.78 5.37$\pm$0.07 8.16$\pm$0.56 65.42$\pm$2.52
Dining room ATISS 37.58$\pm$1.99 137.10$\pm$0.34 8.49$\pm$0.23 23.60$\pm$2.52 67.61$\pm$3.23
DiffuScene 37.87$\pm$2.76 145.48$\pm$1.36 8.63$\pm$0.31 24.08$\pm$1.90 70.57$\pm$2.14
Ours 61.23$\pm$1.67 129.76$\pm$1.61 7.67$\pm$0.18 13.24$\pm$1.79 64.20$\pm$1.90
TABLE II: Quantitive evaluations for zero-shot generative applications on three room types. “Uncond.” stands for unconditional scene synthesis.
Zero-shot Applications Stylization Re-arrangement Completion Uncond.
[…]
Bedroom ATISS 3.44 123.91 61.22 107.67 64.90 89.77 134.51
DiffuScene 1.08 127.35 68.57 106.15 48.57 96.28 135.46
Ours 6.34 122.73 79.59 105.27 69.80 82.98 124.97
Living room ATISS -3.57 110.85 31.97 117.97 43.20 106.48 129.23
DiffuScene -2.69 112.80 41.50 115.30 19.73 95.94 129.75
Ours 0.28 109.39 56.12 106.85 46.94 92.52 117.62
Dining room ATISS -1.11 131.14 36.06 134.54 57.99 122.44 147.52
DiffuScene -2.98 135.20 46.84 133.73 32.34 115.08 150.81
Ours 1.69 128.78 62.08 125.07 60.59 107.86 137.52
4.1 3D Indoor Scene-Instruction Paired Dataset
All experiments are conducted on 3D-FRONT [72], a professionally designed collection of synthetic indoor scenes. However, it does not contain any descriptions of room layouts or object appearances. To construct a high-quality scene-instruction paired dataset, we initially extract view-dependent spatial relations with predefined rules [38; 22]. The dataset is further enhanced by captioning objects with BLIP [27]. To ensure the accuracy of descriptions, the generated captions are filtered by ChatGPT [28; 29] with object ground-truth categories. The final instructions are derived from randomly selected relation triplets. Verbs and conjunctions in the sentences are also randomly picked to keep the diversity and fluency of prompts. Details on 3D dataset curation can be found in Appendix A.1.
4.2 Experiments Settings
4.2.1 Baselines
We compare our method with two state-of-the-art approaches for the 3D scene synthesis tasks:
  • ATISS [17], a Transformer-based auto-regressive network that regards scenes as sets of unordered objects and generates objects and their attributes sequentially.
  • DiffuScene [21], a continuous diffusion model with Gaussian kernels that treats object attributes in one scene as a 2D matrix after padding them to a fixed size.
Both methods can be conditioned on text prompts by cross-attention with a pretrained text encoder. Our preliminary experiments suggest that both baselines encounter difficulties in modeling high-dimensional semantic feature distributions, consequently impacting their performance in generating other attributes. Therefore, we augment them to generate quantized features.
4.2.2 Evaluation Metrics
[…]
4.3 Instruction-drive 3D Scene Synthesis
Table I presents the quantitive evaluations for synthesizing 3D scenes with instructions. We report the average scores of five runs with different random seeds. As demonstrated, even with the enhancement of quantized semantic features, two baseline methods continue to demonstrate inferior performance compared to ours. ATISS outperforms DiffuScene regarding generation fidelity, owing to its capacity to model in discrete spaces. DiffuScene shows better controllability to ATISS because it affords global visibility of samples during generation. Our proposed InstructLayout exhibits the best of both worlds. Remarkably, we achieve a substantial advancement in controllability, measured in iRecall, for scene generative models, surpassing current state-of-the-art approaches by about 15$\sim$25across various room types, all while maintaining high fidelity. It is noteworthy that InstructLayout excels in handling more complex scenes, such as living and dining rooms, which typically comprise an average of 20 objects, in contrast to bedrooms, which have only 8 objects on average, revealing the benefits of modeling intricate 3D scenes associated with the semantic graph prior. Qualitative visualizations are provided in Appendix B.1.
4.4 Zero-shot Applications
Thanks to the discrete design and mask modeling, the learned semantic graph prior can perform diverse downstream tasks without any fine-tuning. We investigate four zero-shot tasks: stylization, re-arrangement, completion, and unconditional generation.
Stylization and re-arrangement task can be formulated as $p_{\phi}(\mathbf{f}|c,\mathbf{s},\mathbf{y})$ and $p_{\phi,\theta}(\mathbf{s}|c,\mathbf{f},\mathbf{y})$ respectively. In the completion task, we intend to add new objects $\{\mathbf{o}_{k}^{i}\}$ to a partial scene $\mathcal{S}_{i}$ with instructions. By filling the partial scene attributes with [MASK] tokens, we treat them as intermediate states during discrete graph denoising, allowing for a straightforward adaptation of the learned semantic graph prior to these tasks in a zero-shot manner. Unconditional synthesis is implemented by simply setting instruction embeddings $\mathbf{y}$ as zeros. To assess controllability in the stylization task, we define $\Delta\coloneqq\frac{1}{N}\sum_{i=1}^{N}\text{CosSim}(\mathbf{f}_{i},\mathbf{d}^{\text{style}}_{i})-\text{CosSim}(\mathbf{f}_{i},\mathbf{d}^{\text{class}}_{i})$, where $\mathbf{d}^{\text{style}}_{i}$ represents the CLIP text feature of object class name with the desired style, and $\mathbf{d}^{\text{class}}_{i}$ is the CLIP text feature with only class information. $\text{CosSim}(\cdot,\cdot)$ calculates the cosine similarity between two vectors.
Evaluations on zero-shot applications are reported in Table II. Our method consistently outperforms two strong baselines in both controllability and fidelity. While ATISS, as an auto-regressive model, is a natural fit for the completion task, its unidirectional dependency chain limits its effectiveness for tasks requiring global scene modeling, such as re-arrangement. DiffuScene can adapt to these tasks by replacing the known parts with the noised corresponding scene attributes during sampling, similar to image in-painting [77; 78]. However, the known attributes are greatly corrupted in the early steps, which could misguide the denoising direction and necessitate fine-tuning. Additionally, DiffuScene also faces challenges in searching for semantic features in a continuous space for stylization. In contrast, InstructLayout globally models scene attributes and treats partial scene attributes as intermediate discrete states during training. These designs effectively eliminate the training-test gap, rendering it highly versatile for a wide range of downstream tasks. Visualizations of zero-shot applications are available in Appendix B.2.
5 InstructLayout for 2D Poster Synthesis
5.1 2D E-commerce Poster-Instruction Paired Dataset
AutoPoster [12] is a recently proposed public poster dataset containing 76,537 E-commerce posters with detailed annotations of layout elements. Compared with an ideal comprehensive poster design pipeline, however, it lacks three essential components, including (1) design instructions describing the desired layout of a poster, (2) principle product descriptions for tagline generation, and (3) clean product images for use as poster backgrounds. To the best of our knowledge, no existing dataset offers comprehensive information for the instruction-driven poster synthesis task [41; 10; 43].
In this work, we bridge this gap by constructing a high-quality 2D E-commerce poster dataset based on the AutoPoster dataset. Similar to 3D scene dataset curation, we use predefined rules as a systematic instruction generation pipeline and collect 73,070 product descriptions for AutoPoster posters from public Internet sources using web crawler techniques. For clean product images, previous works[10; 12] employ an off-the-shelf inpainting model LaMa [79] to remove the graphics elements from the poster. However, LaMa often struggles with posters featuring frequent and significant color shifts, resulting in blurry and unnatural distortion of the processed product images. To mitigate this, we first mask out graphic elements with the average color near the masked region and then apply the LaMa inpainting model to obtain clean product images. More details on the 2D poster dataset curation are provided in Appendix A.2.
TABLE III: Quantitive evaluations for instruction-driven 2D poster synthesis on baselines [6; 7; 10; 19] and InstructLayout. Higher iRecall, Val, Undl and Unds, and lower Ove, N-Ali, and Occ indicate better generation quality. The best results among different models are bolded. Standard deviation values are provided as subscripts.
[…]
LayoutTransformer [6] 65.689$\pm$0.83 98.651$\pm$0.07 0.855$\pm$0.06 0.324$\pm$0.04 95.083$\pm$0.18 84.888$\pm$0.28 7.742$\pm$0.05
LayoutVAE [7] 19.002$\pm$0.57 94.877$\pm$0.10 8.437$\pm$0.06 5.261$\pm$0.10 91.195$\pm$0.25 82.620$\pm$0.35 15.918$\pm$0.03
[…]
DDPM [57] 20.986$\pm$0.50 94.782$\pm$0.23 17.573$\pm$0.45 5.980$\pm$0.21 63.587$\pm$0.56 57.908$\pm$0.56 11.868$\pm$0.20
LayoutDM [19] 62.999$\pm$1.62 93.008$\pm$0.24 2.500$\pm$0.11 0.240$\pm$0.02 86.965$\pm$0.36 82.454$\pm$0.42 7.242$\pm$0.04
Ours 78.451$\pm$0.294 99.977$\pm$0.01 0.785$\pm$0.02 0.165$\pm$0.01 97.123$\pm$0.15 92.490$\pm$0.29 6.565$\pm$0.05
TABLE IV: Quantitive evaluations for generated semantic features on SAP [12] and InstructLayout. “D”, “G”, and “S” represent dominant, gradient, and stroke colors, respectively. Lower MAE and MSE indicate better results.
2D Graphic Features $\downarrow$ D-ab(MSE) $\downarrow$ D-light(MAE) $\downarrow$ S-ab(MSE) $\downarrow$ S-light(MAE) $\downarrow$ G-ab(MSE) $\downarrow$ G-light(MAE)
Random 78.411 5.946 76.062 6.146 80.536 6.596
SAP [12] 74.501 4.197 55.201 2.920 72.369 3.368
Ours 31.047 2.459 33.005 2.872 41.067 1.745
5.2 Experiments Settings
5.2.1 Baselines
We re-implemented several state-of-the-art 2D layout synthesis models for a comprehensive comparison:
  • LayoutTransformer [6], a Transformer decoder that simultaneously predicts categories and discretizes sizes autoregressively.
  • LayoutVAE [7], a VAE-based model that autoregressively predicts the conditional distribution of elements’ counts and bounding boxes.
  • DS-GAN [10], a GAN-based model that builds a Design Sequence GAN and generates layouts with product images in a specially designed order.
  • DDPM [57], a native continuous diffusion model, and we modify it to enable simultaneous prediction of categories and bound boxes.
  • LayoutDM [19], a single-stage discrete diffusion model that discretizes layout attributes and diffuses them in a discrete space.
  • SAP [12], a Transformer encoder designed for graphic semantic feature synthesis with conditions of product images and layout bounding boxes.
During the re-implementation process, we further condition these baselines with product images and text prompts. While DS-GAN [10] fails to converge with textual instructions, we only feed image conditions to it. Note that most of these methods [6; 7; 10; 57; 19] can only generate categories $c$ and spatial features $\mathbf{s}$, and SAP [12] only predict graphic semantic features $\mathbf{f}$.
5.2.2 Evaluation Metrics
Besides iRecall proposed in Sec. 4.2.2 to measure the controllability of layout synthesis, we adopt several graphic and content-aware metrics following previous works [10; 42; 41]. Graphic metrics measure the degree of visual harmony among elements (texts, underlay and logos), including Validity (Val) measuring the validity rate of all predicted elements, Overlay (Ove) representing the average Intersection over Union (IoU) for each pair of elements except underlays, Non-Alignment (N-Ali) measuring the extent of spatial non-alignment, and Underlay-Effectiveness (Undl and Unds) indicating the quality of generated underlay. For the content-aware metric, we use Occlusion (Occ) to evaluate the overall visual impact of the poster by measuring the degree of occlusion of the background subject by elements. For semantic feature quality assessment, we follow Lin et al. [12] to calculate the accuracy of predicted dominant colors, stroke colors, and gradient colors in Lab space using Mean Squared Error (MSE) and Mean Absolute Error (MAE).
5.3 Instruction-drive 2D Poster Synthesis
Table III shows quantitive evaluations on 2D layout generation. We run all models with 5 different random seeds and report the average score. InstructLayout considerably outperforms other existing models by a large margin among all metrics. Especially, our two-stage diffusion model, which skillfully combines discrete and continuous features, remarkably beats both the single-stage diffusion models, the continuous DDPM [57] and the discrete LayoutDM [19], in terms of controllability and visual harmony. It highlights the effectiveness of the proposed discrete graph prior and the capacity of modeling fine spatial relationships facilitated by the continuous diffusion process in the second stage. As presented in Table IV, InstructLayout is also capable of leveraging both image and positional information to mimic the behavior of human designers better and generate graphic semantic features of higher usability. Qualitative results of generated posters with textual content synthesized by the tagline generator are provided in Appendix B.3.
6 Ablation Study
Figure 4: Ablation study on timesteps for diffusion models on 3D scene layout synthesis task.
TABLE V: Ablation study on diffusion timesteps on 2D poster layout synthesis task.
[…]
LayoutDM ($T=110$) 65.841 93.003 2.350 0.222 86.872 82.183 7.236
LayoutDM ($T=60$) 64.293 92.467 2.532 0.279 86.213 81.902 7.293
LayoutDM ($T=35$) 61.825 90.145 2.813 0.468 81.911 75.312 7.470
Ours ($T=100+10$) 78.268 99.995 0.992 0.193 96.067 89.285 6.587
Ours ($T=50+10$) 78.833 99.995 1.151 0.188 95.537 88.560 6.580
Ours ($T=25+10$) 77.862 99.986 1.186 0.192 94.277 86.225 6.662
TABLE VI: Ablation study on different strategies to learn semantic graph prior $p_{\phi}(\mathcal{G}|\mathbf{y})$. “Perm. Invar.” means permutation-invariant graph modeling.
Graph Prior Ours Gaussian Joint Mask Uniform Perm. Invar.
$\uparrow$ iRecall73.64$\pm$1.37 34.18$\pm$2.53 34.21$\pm$2.79 69.22$\pm$3.25 70.49$\pm$2.50
$\downarrow$ FID 114.78$\pm$1.19 128.98$\pm$0.97 130.86$\pm$2.76 139.61$\pm$1.06 116.53$\pm$1.35
[…]
$\downarrow$ KID×${}_{\times\text{1e-3}}$ 0.32$\pm$0.03 2.63$\pm$0.73 4.82$\pm$1.69 10.55$\pm$1.19 0.37$\pm$0.02
     SCA56.02$\pm$0.91 57.10$\pm$3.22 60.37$\pm$3.13 76.79$\pm$3.14 58.64$\pm$1.33
6.1 Diffusion Timesteps
Although containing two diffusion models, our method could achieve better efficiency by reducing the steps of reverse processes without a noticeable decline in performance. We compare InstructLayout with DiffuScene [21] and LayoutDM [19] in Figure 4 and Table V for 3D scene and 2D poster layout synthesis tasks respectively. We try different timesteps $T$ such as “100+100”, “100+10”, and “25+10”, where the first number represents timesteps for semantic graph prior and the latter is for layout decoder. We find that $T=100$ and $10$ is sufficient for $p_{\phi}(\mathcal{G}|\mathbf{y})$ and $p_{\theta}(\mathcal{S}|\mathcal{G})$ respectively. This stems from the fact that each stage in InstructLayout tackles an easier denoising task compared to single-stage diffusion models such as DiffuScene [21] and LayoutDM [11].
6.2 Learning Semantic Graph Prior
We then explore different strategies to learn semantic graph prior. All experiments are conducted on the 3D bedroom dataset. Quantitative results are presented in Table VI.
Transition Matrices for Learning Graph Prior. We investigate the effects of different transition matrices for learning the proposed semantic graph prior, including: (1) Embed all categorical variables into their one-hot encodings and diffuse them by Gaussian kernels, which is similar to Niu et al. [50] and Jo et al. [51]; (2) Jointly masking $\mathbf{F}$ and $\mathbf{E}$ along with nodes $\mathbf{C}$ in a graph, so only the attributes of other objects can be utilized for recovery; (3) Adopt uniform transition matrices without mask states, similar to Vignac et al. [52]. Evaluations on both controllability and fidelity reveal the advantages of our independent mask strategy.
Permutation Non-invariance. Different to previous studies on graph generation [50; 51; 52], we depart from the permutation-invariant modeling convention to ease the semantic graph prior learning process. We strive to preserve exchangeable graph distributions by randomly shuffling object orders during training. Performance for invariant graph prior is provided in the last column of Table VI. Its performance declines due to the unnecessary imposition of invariance in scene synthesis.
7 Conclusion
By integrating a semantic graph prior and a layout decoder, we propose a novel generative framework, InstructLayout, that significantly improves the controllability and fidelity of 2D and 3D layout synthesis, providing a user-friendly interface through instructions in natural languages. Benefiting from the design of semantic graph prior, our method can also apply to diverse applications without any fine-tuning. The controllability and versatility positions InstructLayout as a promising tool. Meanwhile, given the rapid development of large language models (LLMs), integrating an LLM into our instruction-driven pipeline holds significant promise for further enhancing generation interactivity. We hope this work can help in practical scenarios, such as streamlining e-commerce poster design and facilitating interior design, and serves as a good starting point for creating high-quality instruction datasets in 2D poster and 3D scene applications.
Appendix A Dataset Details
A.1 3D Indoor Scene-Instruction Paired Dataset
Following previous works [17; 20; 21], we use three types of indoor rooms in 3D-FRONT [72] and preprocess the dataset by filtering some problematic samples, resulting in 4041 bedrooms, 813 living rooms and 900 dining rooms. The number of objects $N_{i}$ in the valid bedrooms is between 3 and 12 with 21 object categories, i.e., $K_{c}=21$. While for living and dining rooms, $N_{i}$ varies from 3 to 21 and $K_{c}=24$. We use the same data split for training and evaluation as ATISS [17].
The original 3D-FRONT dataset does not contain any descriptions of room layout or object appearance details. To advance research in the field of text-conditional indoor scene generation, we carefully curate a high-quality dataset with paired scenes and instructions for interior design through a multi-step process:
  1. 1.
    Spatial Relation Extraction: View-dependent spatial relations are initially extracted from the 3D-FRONT dataset using predefined rules listed in Appendix A.3.
  2. 2.
    Object Captioning: We further enhance the dataset by providing captions to objects using BLIP [27], a powerful model pretrained for vision-language understanding, given furniture 2D thumbnail images from the original dataset [72].
  3. 3.
    Caption Refinement: As generated captions may not always be accurate, we filter them with corresponding ground-truth categories using ChatGPT [28; 29], a large language model fine-tuned for instruction-based tasks. This results in precise and expressive descriptions of each object in the scene. The prompt and hyperparameters for ChatGPT to filter captions are provided in Appendix A.4.
  4. 4.
    Instruction Generation: The final instructions for scene synthesis are derived from $1\sim 2$ randomly selected “(subject, relation, object)” triplets obtained during the first extraction process. Verbs and conjunctions within sentences are also randomly picked to maintain diversity and fluency.
A.2 2D E-commerce Poster-Instruction Paired Dataset
Several efforts have been made to create public datasets for 2D poster design, as shown in Table VII. CGL-GAN [41] and PosterLayout [10] provide large-scale poster datasets accompanied by bounding box annotations. CapOnImage2M [43] gathers product descriptions and annotated taglines for elements in posters, offering richer information for poster generation. AutoPoster [12] innovatively introduces a dataset featuring tagline and element attribute annotations. However, to the best of our knowledge, no existing dataset provides comprehensive information for instruction-driven poster synthesis.
In this work, we bridge this gap by constructing a high-quality E-commerce poster dataset based on AutoPoster [12]:
  1. 1.
    Instruction Generation: Firstly, we employ predefined rules as shown in Appendix A.3 to extract spatial relations between 2D objects based on the place of their bounding boxes. With the annotation of graphic features, we can derive descriptions of objects in terms of colors, shapes, and fonts. Since the names of colors in LAB space can be ambiguous, we opt to use the nearest match in Web Standard Color palette as the name of each color. Then, we randomly choose several “(subject, relation, object)” triplets and pick up verbs and conjunctions to formulate an instruction.
  2. 2.
    Product Descriptions: We obtained product descriptions from the Internet by web crawler techniques and manually checked and filtered them, resulting in 73,070 pieces of product descriptions.
  3. 3.
    Clean Product Images: In accordance with the improved strategy outlined in Sec. 5.1 for inpainting final posters, we obtained clean product images of significantly higher quality compared to the previous method. The comparisons between previous works [10; 12] are demonstrated in Figure 5.
We provide an example in Figure 6 to showcase various elements contained in our curated poster dataset.
TABLE VII: Comparison of the types of elements contained in various poster datasets.
Dataset Data Type Instruction Product Information Graphic Element Annotation Poster
Image Description Category Bounding Box Attribute Tagline
CGL-GAN [41]
PosterLayout [10]
CapOnImage2M [43]
AutoPoster[12]
Ours
Original
LaMa [79]
Ours
Figure 5: Comparisons of inpainting results between our strategy and the original LaMa [79].
Figure 6: An example from our comprehensive 2D E-commerce poster dataset.
A.3 Relation Definations
Assume $X$ and $Y$ span the ground or poster plane, and $Z$ is the vertical axis for 3D scenes only. We use Center to represent the coordinates of a bounding box’s center. Height is the $Z$-axis size of a 3D bounding box. Relative orientation is computed as $\theta_{so}\coloneqq\text{atan2}(Y_{s}-Y_{o},X_{s}-X_{o})$, where $s$ and $o$ respectively refer to “subject” and “object” in a relationship. $d(s,o)$ is the ground distance between $s$ and $o$. $\text{Inside}(s,o)$ indicates whether the subject center is inside the ground bounding box of the object. We define 11 relationships in both 3D and 2D space as listed in Table IX and Table IX, respectively.
TABLE VIII: Rules to determine the spatial relationships between objects in a 3D scene.
Relationship Rule
Left of ($\theta_{so}\geq\frac{3\pi}{4}$ or $\theta_{so}<-\frac{3\pi}{4}$) and $1<d(s,o)\leq 3$
Right of $-\frac{\pi}{4}\leq\theta_{so}<\frac{\pi}{4}$ and $1<d(s,o)\leq 3$
In front of $\frac{\pi}{4}\leq\theta_{so}<\frac{3\pi}{4}$ and $1<d(s,o)\leq 3$
Behind $-\frac{3\pi}{4}\leq\theta_{so}<-\frac{\pi}{4}$ and $1<d(s,o)\leq 3$
Closely left of ($\theta_{so}\geq\frac{3\pi}{4}$ or $\theta_{so}<-\frac{3\pi}{4}$) and $d(s,o)\leq 1$
Closely right of $-\frac{\pi}{4}\leq\theta_{so}<\frac{\pi}{4}$ and $d(s,o)\leq 1$
Closely in front of $\frac{\pi}{4}\leq\theta_{so}<\frac{3\pi}{4}$ and $d(s,o)\leq 1$
Closely behind $-\frac{3\pi}{4}\leq\theta_{so}<-\frac{\pi}{4}$ and $d(s,o)\leq 1$
Above $(\text{{Center}}_{Z_{s}}-\text{{Center}}_{Z_{o}})>(\text{{Height}}_{s}+\text{{Height}}_{o})/2$
and ($\text{Inside}(s,o)$ or $\text{Inside}(o,s)$)
Below $(\text{{Center}}_{Z_{o}}-\text{{Center}}_{Z_{s}})>(\text{{Height}}_{s}+\text{{Height}}_{o})/2$
and ($\text{Inside}(s,o)$ or $\text{Inside}(o,s)$)
None $d(s,o)>3$
TABLE IX: Rules to determine the spatial relationships between objects in a 2D poster.
Relationship Rule
Left of ($\theta_{so}\geq\frac{3\pi}{4}$ or $\theta_{so}<-\frac{3\pi}{4}$) and $80<d(s,o)\leq 300$
Right of $-\frac{\pi}{4}\leq\theta_{so}<\frac{\pi}{4}$ and $80<d(s,o)\leq 300$
above $\frac{\pi}{4}\leq\theta_{so}<\frac{3\pi}{4}$ and $80<d(s,o)\leq 300$
below $-\frac{3\pi}{4}\leq\theta_{so}<-\frac{\pi}{4}$ and $80<d(s,o)\leq 300$
Closely left of ($\theta_{so}\geq\frac{3\pi}{4}$ or $\theta_{so}<-\frac{3\pi}{4}$) and $d(s,o)\leq 80$
Closely right of $-\frac{\pi}{4}\leq\theta_{so}<\frac{\pi}{4}$ and $d(s,o)\leq 80$
Closely above $\frac{\pi}{4}\leq\theta_{so}<\frac{3\pi}{4}$ and $d(s,o)\leq 80$
Closely below $-\frac{3\pi}{4}\leq\theta_{so}<-\frac{\pi}{4}$ and $d(s,o)\leq 80$
Enclosing Inside($o$,$s$)
Enclosed with Inside($s$,$o$)
None $d(s,o)>300$
A.4 Caption Refinement by ChatGPT
The generated 3D object captions from BLIP are refined by ChatGPT (gpt-3.5-turbo). Our prompt to ChatGPT is provided in Table X. We set the hyperparameter temperature and top_p for text generation to 0.2 and 0.1 respectively, encouraging more deterministic and focused outputs.
TABLE X: Prompt for ChatGPT to refine raw 3D object descriptions.
Given a description of furniture from a captioning model and its ground-truth category, please combine their information and generate a new short description in one line. The provided category must be the descriptive subject of the new description. The new description should be as short and concise as possible, encoded in ASCII. Do not describe the background and counting numbers. Do not describe size like ‘small’, ‘large’, etc. Do not include descriptions like ‘a 3D model’, ‘a 3D image’, ‘a 3D printed’, etc. Descriptions such as color, shape and material are very important, you should include them. If the old description is already good enough, you can just copy it. If the old description is meaningless, you can just only include the category. For example: Given ‘a 3D image of a brown sofa with four wooden legs’ and ‘multi-seat sofa’, you should return: a brown multi-seat sofa with wooden legs. Given ‘a pendant lamp with six hanging balls on the white background’ and ‘pendant lamp’, you should return: a pendant lamp with hanging balls. Given ‘a black and brown chair with a floral pattern’ and ‘armchair’, you should return: a black and brown floral armchair. The above examples indicate that you should delete the redundant words in the old description, such as ‘3D image’, ‘four’, ‘six’ and ‘white background’, and you must include the category name as the subject in the new description. The old descriptions is ‘{BLIP caption}’, its category is ‘{ground-truth category}’, the new descriptions should be:
Appendix B Qualitative Results
B.1 Instruction-driven 3D Scene Synthesis
We present visualizations of instruction-driven synthesized 3D scenes in Figure 7, 8 and 9. Besides the quantitative evaluations shown in Table I, these qualitative visualizations also show the superiority of our method over previous state-of-the-art approaches in terms of adherence to instructions and generative quality.
(a) Instructions
(b) ATISS
(c) DiffuScene
(d) Ours
Figure 7: Visualizations for instruction-drive synthesized 3D bedrooms by ATISS [17], DiffuScene [21] and our method.
(a) Instructions
(b) ATISS
(c) DiffuScene
(d) Ours
Figure 8: Visualizations for instruction-drive synthesized 3D living rooms by ATISS [17], DiffuScene [21] and our method.
(a) Instructions
(b) ATISS
(c) DiffuScene
(d) Ours
Figure 9: Visualizations for instruction-drive synthesized 3D dining rooms by ATISS [17], DiffuScene [21] and our method.
B.2 Zero-shot 3D Scene Applications
We present visualizations illustrating various zero-shot instruction-driven applications, including stylization, re-arrangement, completion, and unconditional 3D scene synthesis in Figure 10, 11, 12 and 13 respectively. We find that the autoregressive model ATISS [17] tends to generate redundant objects, resulting in chaotic synthesized scenes. DiffuScene [21] encounters challenges in accurately modeling object semantic features, often yielding objects that lack coherence in terms of style or pairing, thereby diminishing the aesthetic appeal of the synthesized scenes. Moreover, both baseline models frequently struggle to follow the provided instructions during conditional generation.
(a) Instructions
(b) Original Scenes
(c) ATISS
(d) DiffuScene
(e) Ours
Figure 10: Visualizations for instruction-drive 3D scenes stylization by ATISS [17], DiffuScene [21] and our method.
(a) Instructions
(b) Messy Scenes
(c) ATISS
(d) DiffuScene
(e) Ours
Figure 11: Visualizations for instruction-drive 3D scenes re-arrangement by ATISS [17], DiffuScene [21] and our method.
(a) Instructions
(b) Partial Scenes
(c) ATISS
(d) DiffuScene
(e) Ours
Figure 12: Visualizations for instruction-drive 3D scenes completion by ATISS [17], DiffuScene [21] and our method.
(a) ATISS
(b) DiffuScene
(c) Ours
Figure 13: Visualizations for unconditional 3D scenes stylization by ATISS [17], DiffuScene [21] and our method.
B.3 Instruction-driven 2D Poster Synthesis
Comparisons of our synthesized 2D posters with other state-of-the-art methods are shown in Figure 14, 15, and 16, demonstrating the superiority of our method compared to previous state-of-the-art approaches in terms of better mimicking the behavior of human designers and following provided instructions.
(a) Instruction & Product Information
(b) LayoutTrans
(c) LayoutVAE
(d) LayoutDM
(e) Ours
(f) Ours (Rendered)
Instruction
Information
Synthesized Taglines: [Children’s Warmth Essential; Stylish Jacket for Cozy Growth.]
Instruction
Information
Synthesized Taglines: [Vietnamese Pure Sand Gold; Customized Happiness.]
Instruction
Information
Synthesized Taglines: [Raccoon Fur Down Jacket; Winter Essential;
Brave the Cold with Ease, Lightweight and Comfortable.]
Instruction
Information
Synthesized Taglines: [Vintage Purple Clay Tea Cup; Japanese Fine Carving; Custom Logo Design.]
Figure 14: Visualizations for instruction-drive synthesized 2D posters by LayoutTrans [6], LayoutVAE [7], LayoutDM [19] and our method. Note that posters are rendered according to synthesized graphic and textual features.
(a) Instruction & Product Information
(b) LayoutTrans
(c) LayoutVAE
(d) LayoutDM
(e) Ours
(f) Ours (Rendered)
Instruction
Information
Synthesized Taglines: [European Fashion White Shoes; Internal Height Increase Design.]
Instruction
Information
Synthesized Taglines: [Deer Crystal Porcelain Wall Decoration for the Foyer; Light Luxury Modern Minimalist Decor.]
Instruction
Information
Synthesized Taglines: [Wild&Wolf Coffee Cup; Delicate Enamel Hand Brewing for Outdoor Use; British Style.]
Instruction
Information
Synthesized Taglines: [KeFiaYa Jacquard Half Skirt; A-line Style Short Skirt for Petite and Fashionable Looks.]
Figure 15: Visualizations for instruction-drive synthesized 2D posters by LayoutTrans [6], LayoutVAE [7], LayoutDM [19] and our method. Note that posters are rendered according to synthesized graphic and textual features.
(a) Instruction & Product Information
(b) LayoutTrans
(c) LayoutVAE
(d) LayoutDM
(e) Ours
(f) Ours (Rendered)
Instruction
Information
Synthesized Taglines: [Nordic Style Technology Fabric Sofa; Light Luxury Modern Living Room Centerpiece.]
Instruction
Information
Synthesized Taglines: [Smart Neck Massage; Voice-Controlled One-Button Neck Protection.]
Instruction
Information
Synthesized Taglines: [Password Suitcase; Personalized Graffiti Fashion Trend; Violent Bear Luggage.]
Instruction
Information
Synthesized Taglines: [Autumn and winter essential; Fashionable fleece-lined denim jacket;
Loose design for warmth and thickness; 2023 New Style.]
Figure 16: Visualizations for instruction-drive synthesized 2D posters by LayoutTrans [6], LayoutVAE [7], LayoutDM [19] and our method. Note that posters are rendered according to synthesized graphic and textual features.

### v3

@@
 1 Introduction
-[…]
-[lines=2]Automatically synthesizing controllable and appealing 2D and 3D layouts has been a challenging task for computer vision and graphics [1; 2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12; 13; 14; 15], which requires building a generative model by multimodal conditional information.
+Automatically synthesizing controllable and appealing 2D and 3D layouts has been a challenging task for computer vision and graphics [1; 2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12; 13; 14], which requires building a generative model by multimodal conditional information.
 An ideal layout synthesis system should fulfill at least three objectives: (1) comprehending instructions in natural languages, thus providing an intuitive and user-friendly interface; (2) designing object compositions that exhibit aesthetic appeal and thematic harmony; (3) placing objects in appropriate positions adhering to their functions and regular arrangements.
@@
 However, natural instructions for layout design often rely on abstract object relationships, which has posed a significant challenge for recent advancements in 2D and 3D layout synthesis.
-Previous works [16; 17; 18; 19; 20; 21] primarily exclude instructional prompts or do not prioritize them, leading to a low level of controllability and interpretability in the generated layouts.
-Other studies [22; 23; 24] utilize relation graphs to provide explicit control over object interactions, which are, however, too complicated and fussy for human users to specify.
-Moreover, previous works merely represent objects by categories  [22; 6; 19; 17] or low-dimensional features [25; 11; 12; 21] which lack visual appearance details, resulting in style inconsistency and constraining customization options in layout synthesis.
+Previous works [15; 16; 17; 18; 19; 20] primarily exclude instructional prompts or do not prioritize them, leading to a low level of controllability and interpretability in the generated layouts.
+Other studies [21; 22; 23] utilize relation graphs to provide explicit control over object interactions, which are, however, too complicated and fussy for human users to specify.
+Moreover, previous works merely represent objects by categories  [21; 4; 18; 16] or low-dimensional features [24; 9; 10; 20] which lack visual appearance details, resulting in style inconsistency and constraining customization options in layout synthesis.
 To address these issues, we present InstructLayout, a novel generative framework for both 2D and 3D layout synthesis with natural language instructions.
@@
 Consequently, our model can handle these discrete features in a unified manner.
-To fit practical scenarios and promote the benchmarking of instruction-drive layout synthesis, we curate two high-quality datasets containing paired layouts and instructions with the help of large language and multimodal models [27; 28; 29] for both 2D E-commerce poster and 3D indoor scene synthesis tasks.
+To fit practical scenarios and promote the benchmarking of instruction-drive layout synthesis, we curate two high-quality datasets containing paired layouts and instructions with the help of large language and multimodal models [26; 27; 28] for both 2D E-commerce poster and 3D indoor scene synthesis tasks.
 Comprehensive quantitative evaluations reveal that InstructLayout surpasses previous state-of-the-art methods by a large margin in terms of both generation controllability and fidelity.
@@
 2.1 3D Scene Layout Synthesis
-Graphs have been used to guide complex scene synthesis in the form of scene hierarchies [30; 31], parse trees [32], scene graphs [33; 34], etc.
+Graphs have been used to guide complex scene synthesis in the form of scene hierarchies [29; 30], parse trees [31], scene graphs [32; 33], etc.
 Wang et al.
@@
 They also adopt an autoregressive model [34] to generate relation graphs, which are unconditional and have limited object attributes.
-Other works [22; 23; 24] adopt conditional VAEs [36; 37] with graph convolutional networks [38] to generate layouts.
+Other works [21; 22; 23] adopt conditional VAEs [35; 36] with graph convolutional networks [37] to generate layouts.
 While offering high controllability, these methods demand the specification of elaborate graph conditions, which are notably more intricate than those driven by natural languages.
-With the advent of attention mechanisms [39], recent approaches [16; 17; 20; 21] can implicitly acquire object relations by self-attention and condition scene synthesis with texts by cross-attention.
+With the advent of attention mechanisms [38], recent approaches [15; 16; 19; 20] can implicitly acquire object relations by self-attention and condition scene synthesis with texts by cross-attention.
 However, text prompts in these works tend to be relatively simple, containing only object categories or lacking layout descriptions, limiting the expressiveness and customization.
@@
 Numerous methods have been proposed to automatically synthesize 2D layouts for posters, user interfaces, E-commerce advertisements, etc.
-Prior research [7; 6; 40; 11; 19] has primarily concentrated on generating layouts from scratch on a blank canvas, while others [18; 41; 42; 10; 12] focus on image content-aware layout synthesis, leading to more practical applications.
-However, they are either confined to unconditional generation based solely on the canvas(blank or image prior) [7; 6; 41; 42] or relying on object categories [11; 19], lacking the capability to control the layout synthesis process using user-friendly instructions.
-Furthermore, due to the scarcity of high-quality datasets, only a few studies [12; 43] have delved into predicting 2D object attributes and their textual content, which is vital for generating realistic and informative layouts.
+Prior research [5; 4; 39; 9; 18] has primarily concentrated on generating layouts from scratch on a blank canvas, while others [17; 40; 41; 8; 10] focus on image content-aware layout synthesis, leading to more practical applications.
+However, they are either confined to unconditional generation based solely on the canvas(blank or image prior) [5; 4; 40; 41] or relying on object categories [9; 18], lacking the capability to control the layout synthesis process using user-friendly instructions.
+Furthermore, due to the scarcity of high-quality datasets, only a few studies [10; 42] have delved into predicting 2D object attributes and their textual content, which is vital for generating realistic and informative layouts.
 2.3 Generative Models for Graphs
-There have been lots of endeavors on generative models for undirected graphs, molecules and scene graphs by autoregressive models [44; 45], VAEs [46; 47], GANs [48; 49] and diffusion models [50; 51; 52; 53].
+There have been lots of endeavors on generative models for undirected graphs, molecules and scene graphs by autoregressive models [43; 44], VAEs [45; 46], GANs [47; 48] and diffusion models [49; 50; 51; 52].
 However, none of them is capable of understanding text conditions.
@@
 Given instructions $\mathbf{y}$, our goal is to learn the conditional scene distribution $q(\mathcal{S}|\mathbf{y})$.
-Rather than direct modeling [17; 19; 12; 21], we employ well-structured and informative graphs to serve as general and semantic latent.
+Rather than direct modeling [16; 18; 10; 20], we employ well-structured and informative graphs to serve as general and semantic latent.
 Each graph $\mathcal{G}_{i}$ contains a node set $\mathcal{V}_{i}\coloneqq\{\mathbf{v}_{j}^{i}\}_{j=1}^{N_{i}}$ and a directed edge set $\mathcal{E}_{i}\coloneqq\{e_{jk}^{i}|\mathbf{v}_{j}^{i},\mathbf{v}_{k}^{i}\in\mathcal{V}_{i}\}$.
@@
 Diffusion models can accommodate both continuous and categorical distributions by a meticulous design of the diffusion kernel $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$.
-A prominent choice for continuous variables is the Gaussian kernel [57; 58; 21].
-Uniform transition is widely adopted for categorical variables [59; 60; 52].
+A prominent choice for continuous variables is the Gaussian kernel [56; 57; 20].
+Uniform transition is widely adopted for categorical variables [58; 59; 51].
 In this work, we utilize both the discrete and continuous diffusion models for the two stages in InstructLayout respectively to model distinct layout attributes.
@@
 However, these high-dimensional features (e.g., $d=1280$ in OpenShape pointbert-vitg14) are too complicated to model.
-We circumvent this drawback by introducing a vector-quantized variational autoencoder (VQ-VAE) [62; 63] for feature vectors and quantize each of them by several learnable tokens from the maintained codebook, as depicted in Figure 2(a).
+We circumvent this drawback by introducing a vector-quantized variational autoencoder (VQ-VAE) [61; 62] for feature vectors and quantize each of them by several learnable tokens from the maintained codebook, as depicted in Figure 2(a).
 The intuition behind it is that 3D models share general intrinsic characteristics, encompassing attributes like colors, materials, and basic geometric shapes.
@@
 3.3.2 Discrete Semantic Graph Diffusion
-After feature discretization, all attributes in a semantic graph are categorical: a graph $\mathcal{G}_{i}\coloneqq(\mathcal{C}_{i},\mathcal{F}_{i},\mathcal{E}_{i})$ with $N_{i}$ nodes, where $\mathcal{C}_{i}\coloneqq\{1,\dots,K_{c}\}^{N_{i}}$, $\mathcal{E}_{i}\coloneqq\{1,\dots,K_{e}\}^{N_{i}\times N_{i}}$ and $\mathcal{F}_{i}\coloneqq\{1,\dots,K_{f}\}^{N_{i}\times n_{f}}$, where $n_{f}$ is the number of discrete features.
-While it is possible to embed discrete variables in continuous spaces using one-hot encodings, it diminishes the sparsity inherent in the original data and imposes a substantial burden on network optimization.
-Instead, we propose to model the semantic graph prior by a mask-based discrete diffusion model.
-For a scalar discrete random variable with $K$ categories $x\in\{1,\dots,K\}$, diffusion noise is defined by a series of transition matrices $\mathbf{Q}\in\mathbb{R}^{K\times K}$.
-The forward process at timestep $t$ is expressed as $q(x_{t}|x_{t-1})\coloneqq\mathbf{x}_{t}^{\top}\mathbf{Q}_{t}\mathbf{x}_{t-1}$, where $\mathbf{x}_{t}\in\mathbb{R}^{K}$ is the column one-hot encoding for $x_{t}$ and $[\mathbf{Q}_{t}]_{mn}\coloneqq q(x_{t}=m|x_{t-1}=n)$ is the probability that $x_{t-1}$ transits to the category $m$ from $n$.
+After feature discretization, all attributes in a semantic graph $\mathcal{G}_{i}\coloneqq(\mathcal{C}_{i},\mathcal{F}_{i},\mathcal{E}_{i})$ are categorical, where $\mathcal{C}_{i}\in\{1,\dots,K_{c}\}^{N_{i}}$, $\mathcal{F}_{i}\in\{1,\dots,K_{f}\}^{N_{i}\times n_{f}}$, and $\mathcal{E}_{i}\in\{1,\dots,K_{e}\}^{N_{i}\times N_{i}}$, where $n_{f}$ is the number of discrete features.
+While continuous embeddings such as one-hot vectors are commonly used, they often obscure the inherent sparsity of categorical data and complicate optimization.
+We instead model the semantic graph prior using a mask-based discrete diffusion process.
+For a categorical variable $x\in\{1,\dots,K\}$, diffusion noise is defined by transition matrices $\mathbf{Q}t\in\mathbb{R}^{K\times K}$ over timesteps $t$.
+The forward process is given by $q(x_{t}|x_{t-1})\coloneqq\mathbf{x}_{t}^{\top}\mathbf{Q}_{t}\mathbf{x}_{t-1}$, where $\mathbf{x}_{t}$ is the column one-hot encoding of $x_{t}$ and $[\mathbf{Q}_{t}]_{mn}\coloneqq q(x_{t}=m|x_{t-1}=n)$ is the probability that $x_{t-1}$ transits to the category $m$ from $n$.
 The probabilistic distribution of $x_{t}$ can be directly derived from $x_{0}$: $q(x_{t}|x_{0})\coloneqq\mathbf{x}_{t}^{\top}\bar{\mathbf{Q}}_{t}\mathbf{x}_{0}$, where $\bar{\mathbf{Q}}_{t}\coloneqq\mathbf{Q}_{t}\cdots\mathbf{Q}_{1}$.
-Instead of commonly used Gaussian or uniform transitions for graph generation [50; 59; 51; 52], we propose to diffuse semantic graphs by independently masking graph attributes (i.e., object class $c$, discrete feature indices $f$ and relation $e$) by introducing an absorbing state [MASK] [60; 64] to each uniform transition matrix.
-For object class $c$, its transition matrix is defined as:
+Instead of relying on commonly used Gaussian or uniform transitions for graph generation [49; 58; 50; 51], we adopt a diffusion process tailored for semantic graphs by independently masking graph attributes (i.e., object class $c$, discrete feature indices $f$, and relation $e$) by introducing an absorbing state, [MASK] [59; 63], into each uniform transition matrix.
+For the object class $c$, its transition matrix is defined as:
 $\mathbf{Q}_{t}^{\mathbf{C}}\coloneqq\left[\begin{array}[]{ccccc}\alpha_{t}^{c}+\beta_{t}^{c}&\beta_{t}^{c}&\cdots&\beta_{t}^{c}&0\\ \beta_{t}^{c}&\alpha_{t}^{c}+\beta_{t}^{c}&\cdots&\beta_{t}^{c}&0\\ \vdots&\vdots&\ddots&\beta_{t}^{c}&0\\ \beta_{t}^{c}&\beta_{t}^{c}&\beta_{t}^{c}&\alpha_{t}^{c}+\beta_{t}^{c}&0\\ \gamma_{t}^{c}&\gamma_{t}^{c}&\gamma_{t}^{c}&\gamma_{t}^{c}&1\end{array}\right],$ (3)
-by which $c_{t}$ has a probability of $\gamma_{t}^{c}$ to be masked, a probability of $\alpha_{t}^{c}$ to maintain the same, leaving a chance of $1-\gamma_{t}^{c}-\alpha_{t}^{c}$ for uniform sampling.
-[MASK] will always stay in its state.
-Transition matrices for $f$ and $e$, denoted as $\mathbf{Q}_{t}^{\mathbf{F}}$ and $\mathbf{Q}_{t}^{\mathbf{E}}$ respectively, exhibit analogous structures.
-Schedules of $(\alpha_{t},\beta_{t},\gamma_{t})$ are designed to admit that the initial states for semantic graphs are all masked.
+by which $c_{t}$ has a probability of $\gamma_{t}^{c}$ to be masked, a probability of $\alpha_{t}^{c}$ to remain unchanged, and a remaining probability of $1-\gamma_{t}^{c}-\alpha_{t}^{c}$ to be uniformly resampled from the label space.
+The absorbing state [MASK] remains fixed once reached.
+Transition matrices for $f$ and $e$, denoted by $\mathbf{Q}_{t}^{\mathbf{F}}$ and $\mathbf{Q}_{t}^{\mathbf{E}}$ respectively, follow the same design.
+The schedules of $(\alpha_{t},\beta_{t},\gamma_{t})$ are crafted such that semantic graphs start fully masked at the beginning of the diffusion process.
 Since the number of objects varies across different layouts, semantic graphs are padded by empty states to maintain a consistent number of $N$ objects.
 One-hot encodings for scalar variables $c$, $f$ and $e$ in a layout are denoted as $\mathbf{C}\in\mathbb{R}^{N\times(K_{c}+2)}$, $\mathbf{F}\in\mathbb{R}^{N\times n_{f}\times(K_{f}+2)}$ and $\mathbf{E}\in\mathbb{R}^{N\times N\times(K_{e}+2)}$ respectively.
-Here, “$+2$” accounts for each variable’s two extra states (i.e., empty state and mask state).
+Here, “+2+2+ 2” accounts for each variable’s two extra states (i.e., empty state and mask state).
 A one-hot encoded semantic graph $G_{0}\coloneqq(\mathbf{C}_{0},\mathbf{F}_{0},\mathbf{E}_{0})$ at timestep $t$ is formulated as
@@
   • The introduction of mask states facilitates the distinction between corrupted and clean states, thus simplifying the denoising task.
-These benefits are critical, especially for intricate semantic graphs and diverse downstream generative tasks, compared with simple graph generative tasks [50; 51; 52].
+These benefits are critical, especially for intricate semantic graphs and diverse downstream generative tasks, compared with simple graph generative tasks [49; 50; 51].
 An ablation study on the choice of $\mathbf{Q}$ is provided in Sec.
@@
 For the 2D layout synthesis task, although spatial relations can be inferred from edges in the semantic graph, we additionally provide clean product poster images $\mathbf{y}_{\text{image}}$ as conditions for more precise arrangements and to avoid occluding the principal products.
-A continuous diffusion model with variance-preserving Gaussian kernels [57; 58] is adopted to learn $p_{\theta}(\mathbf{L}|\mathcal{G})$.
+A continuous diffusion model with variance-preserving Gaussian kernels [56; 57] is adopted to learn $p_{\theta}(\mathbf{L}|\mathcal{G})$.
 Following Ho et al.
@@
 3.4.2 Tagline Generation for 2D Poster
-After synthesizing semantic features $\mathbf{f}$ and spatial features $\mathbf{s}$, the pipeline for 3D scene synthesis is completed by retrieving 3D objects from an off-the-shelf dataset based on the generated $\mathbf{f}$ and transforming them to appropriate positions using the generated $\mathbf{s}$.
-For 2D poster synthesis, we proceed to predict the textual contents in bounding boxes using all generated $\mathbf{f}$ and $\mathbf{s}$ associated with user-provided product descriptions and images.
-Concretely, we primarily adopt the method proposed by Gao et al.
-[43] to generate tagline captions.
-This method involves an autoregressive transformer decoder considering multimodal conditions such as background images, product descriptions, and spatial features.
-A progressive training strategy trains it through two tasks, including caption generation and caption matching.
+After synthesizing semantic features $\mathbf{f}$ and spatial layouts $\mathbf{s}$, the pipeline for 3D scene synthesis is completed by retrieving 3D objects from an off-the-shelf dataset based on the generated $\mathbf{f}$ and transforming them to appropriate positions using the generated $\mathbf{s}$.
+For 2D poster synthesis, we further generate textual taglines for each bounding box using all generated $\mathbf{f}$, $\mathbf{s}$ and user-provided product descriptions and images.
+We adopt the multimodal captioning framework proposed by Gao et al.
+[42], which uses an autoregressive Transformer decoder conditioned on background images, product descriptions and spatial layout.
+The model is trained progressively via caption generation and matching tasks.
 However, the absence of clean product images and semantic features of text bounding boxes results in suboptimal tagline generation.
 To address this issue, we utilize high-quality cleaned product images processed according to the strategy proposed in Sec.
-5.1 and incorporate the generated semantic features from Sec 3.3 as an additional robust condition to guide generation.
+5.1 and incorporate the generated semantic features from Sec 3.3 as an additional conditioning.
+This results in more coherent and product-relevant tagline generation.
 3.5 Model Architecture
@@
 Object categories and features together form input tokens for Transformers.
-Message passing on graphs is operated via node self-attention and node-edge fusion with FiLM [66], which linearly modulates edge embeddings and node attention matrices before softmax [67; 52].
-Timestep for diffusion $t$ is injected by AdaLN [68; 69].
+Message passing on graphs is operated via node self-attention and node-edge fusion with FiLM [65], which linearly modulates edge embeddings and node attention matrices before softmax [66; 51].
+Timestep for diffusion $t$ is injected by AdaLN [67; 68].
 In the graph prior $p_{\phi}(\mathcal{G}|\mathbf{y})$, instructions and optional product images are embedded by a frozen multimodal encoder CLIP [69] and consistently influence network outputs by cross-attention mechanisms.
@@
 Although $\mathcal{G}$ should ideally remain invariant to node permutations, invariant diffusion models could encounter learning challenges for multi-mode modeling.
-Thus, each node feature is added with positional encodings [39; 71; 21] before the permutation-equivariant Transformer.
+Thus, each node feature is added with positional encodings [38; 70; 20] before the permutation-equivariant Transformer.
 Exchangeability for graph prior distributions is strived by random permutation augmentation during the training process.
@@
 TABLE II: Quantitive evaluations for zero-shot generative applications on three room types. “Uncond.” stands for unconditional scene synthesis.
-Zero-shot Applications Stylization Re-arrangement Completion Uncond.
+Zero-shot
+Applications
+Stylization Re-arrangement Completion Uncond.
 […]
@@
 However, it does not contain any descriptions of room layouts or object appearances.
-To construct a high-quality scene-instruction paired dataset, we initially extract view-dependent spatial relations with predefined rules [38; 22].
+To construct a high-quality scene-instruction paired dataset, we initially extract view-dependent spatial relations with predefined rules [37; 21].
 The dataset is further enhanced by captioning objects with BLIP [26].
-To ensure the accuracy of descriptions, the generated captions are filtered by ChatGPT [28; 29] with object ground-truth categories.
+To ensure the accuracy of descriptions, the generated captions are filtered by ChatGPT [27; 28] with object ground-truth categories.
 The final instructions are derived from randomly selected relation triplets.
@@
 It is noteworthy that InstructLayout excels in handling more complex scenes, such as living and dining rooms, which typically comprise an average of 20 objects, in contrast to bedrooms, which have only 8 objects on average, revealing the benefits of modeling intricate 3D scenes associated with the semantic graph prior.
-Qualitative visualizations are provided in Appendix B.1.
+Qualitative visualizations are provided in Appendix D.1.
 4.4 Zero-shot Applications
@@
 We investigate four zero-shot tasks: stylization, re-arrangement, completion, and unconditional generation.
-Stylization and re-arrangement task can be formulated as $p_{\phi}(\mathbf{f}|c,\mathbf{s},\mathbf{y})$ and $p_{\phi,\theta}(\mathbf{s}|c,\mathbf{f},\mathbf{y})$ respectively.
-In the completion task, we intend to add new objects $\{\mathbf{o}_{k}^{i}\}$ to a partial scene $\mathcal{S}_{i}$ with instructions.
-By filling the partial scene attributes with [MASK] tokens, we treat them as intermediate states during discrete graph denoising, allowing for a straightforward adaptation of the learned semantic graph prior to these tasks in a zero-shot manner.
-Unconditional synthesis is implemented by simply setting instruction embeddings $\mathbf{y}$ as zeros.
-To assess controllability in the stylization task, we define $\Delta\coloneqq\frac{1}{N}\sum_{i=1}^{N}\text{CosSim}(\mathbf{f}_{i},\mathbf{d}^{\text{style}}_{i})-\text{CosSim}(\mathbf{f}_{i},\mathbf{d}^{\text{class}}_{i})$, where $\mathbf{d}^{\text{style}}_{i}$ represents the CLIP text feature of object class name with the desired style, and $\mathbf{d}^{\text{class}}_{i}$ is the CLIP text feature with only class information.
-$\text{CosSim}(\cdot,\cdot)$ calculates the cosine similarity between two vectors.
+[…]
 Evaluations on zero-shot applications are reported in Table II.
@@
 While ATISS, as an auto-regressive model, is a natural fit for the completion task, its unidirectional dependency chain limits its effectiveness for tasks requiring global scene modeling, such as re-arrangement.
-DiffuScene can adapt to these tasks by replacing the known parts with the noised corresponding scene attributes during sampling, similar to image in-painting [77; 78].
+DiffuScene can adapt to these tasks by replacing the known parts with the noised corresponding scene attributes during sampling, similar to image in-painting [76; 77].
 However, the known attributes are greatly corrupted in the early steps, which could misguide the denoising direction and necessitate fine-tuning.
@@
 These designs effectively eliminate the training-test gap, rendering it highly versatile for a wide range of downstream tasks.
-Visualizations of zero-shot applications are available in Appendix B.2.
+Visualizations of zero-shot applications are available in Appendix D.2.
 5 InstructLayout for 2D Poster Synthesis
@@
 Compared with an ideal comprehensive poster design pipeline, however, it lacks three essential components, including (1) design instructions describing the desired layout of a poster, (2) principle product descriptions for tagline generation, and (3) clean product images for use as poster backgrounds.
-To the best of our knowledge, no existing dataset offers comprehensive information for the instruction-driven poster synthesis task [41; 10; 43].
+To the best of our knowledge, no existing dataset offers comprehensive information for the instruction-driven poster synthesis task [40; 8; 42].
 In this work, we bridge this gap by constructing a high-quality 2D E-commerce poster dataset based on the AutoPoster dataset.
 Similar to 3D scene dataset curation, we use predefined rules as a systematic instruction generation pipeline and collect 73,070 product descriptions for AutoPoster posters from public Internet sources using web crawler techniques.
-For clean product images, previous works[10; 12] employ an off-the-shelf inpainting model LaMa [79] to remove the graphics elements from the poster.
+For clean product images, previous works[8; 10] employ an off-the-shelf inpainting model LaMa [78] to remove the graphics elements from the poster.
 However, LaMa often struggles with posters featuring frequent and significant color shifts, resulting in blurry and unnatural distortion of the processed product images.
@@
 More details on the 2D poster dataset curation are provided in Appendix A.2.
-TABLE III: Quantitive evaluations for instruction-driven 2D poster synthesis on baselines [6; 7; 10; 19] and InstructLayout.
+TABLE III: Quantitive evaluations for instruction-driven 2D poster synthesis on baselines [4; 5; 8; 18] and InstructLayout.
 Higher iRecall, Val, Undl and Unds, and lower Ove, N-Ali, and Occ indicate better generation quality.
@@
 While DS-GAN [8] fails to converge with textual instructions, we only feed image conditions to it.
-Note that most of these methods [6; 7; 10; 57; 19] can only generate categories $c$ and spatial features $\mathbf{s}$, and SAP [12] only predict graphic semantic features $\mathbf{f}$.
+Note that most of these methods [4; 5; 8; 56; 18] can only generate categories $c$ and spatial features $\mathbf{s}$, and SAP [10] only predict graphic semantic features $\mathbf{f}$.
 5.2.2 Evaluation Metrics
 Besides iRecall proposed in Sec.
-4.2.2 to measure the controllability of layout synthesis, we adopt several graphic and content-aware metrics following previous works [10; 42; 41].
+4.2.2 to measure the controllability of layout synthesis, we adopt several graphic and content-aware metrics following previous works [8; 41; 40].
 Graphic metrics measure the degree of visual harmony among elements (texts, underlay and logos), including Validity (Val) measuring the validity rate of all predicted elements, Overlay (Ove) representing the average Intersection over Union (IoU) for each pair of elements except underlays, Non-Alignment (N-Ali) measuring the extent of spatial non-alignment, and Underlay-Effectiveness (Undl and Unds) indicating the quality of generated underlay.
@@
 As presented in Table IV, InstructLayout is also capable of leveraging both image and positional information to mimic the behavior of human designers better and generate graphic semantic features of higher usability.
-Qualitative results of generated posters with textual content synthesized by the tagline generator are provided in Appendix B.3.
+Qualitative results of generated posters with textual content synthesized by the tagline generator are provided in Appendix D.3.
 6 Ablation Study
@@
 […]
-$\downarrow$ KID×${}_{\times\text{1e-3}}$ 0.32$\pm$0.03 2.63$\pm$0.73 4.82$\pm$1.69 10.55$\pm$1.19 0.37$\pm$0.02
+[…]
    SCA56.02$\pm$0.91 57.10$\pm$3.22 60.37$\pm$3.13 76.79$\pm$3.14 58.64$\pm$1.33
@@
 We try different timesteps $T$ such as “100+100”, “100+10”, and “25+10”, where the first number represents timesteps for semantic graph prior and the latter is for layout decoder.
-We find that $T=100$ and $10$ is sufficient for $p_{\phi}(\mathcal{G}|\mathbf{y})$ and $p_{\theta}(\mathcal{S}|\mathcal{G})$ respectively.
+We find that $T=100$ and 101010 is sufficient for $p_{\phi}(\mathcal{G}|\mathbf{y})$ and $p_{\theta}(\mathcal{S}|\mathcal{G})$ respectively.
 This stems from the fact that each stage in InstructLayout tackles an easier denoising task compared to single-stage diffusion models such as DiffuScene [20] and LayoutDM [9].
@@
 Permutation Non-invariance.
-Different to previous studies on graph generation [50; 51; 52], we depart from the permutation-invariant modeling convention to ease the semantic graph prior learning process.
+Different to previous studies on graph generation [49; 50; 51], we depart from the permutation-invariant modeling convention to ease the semantic graph prior learning process.
 We strive to preserve exchangeable graph distributions by randomly shuffling object orders during training.
@@
 A.1 3D Indoor Scene-Instruction Paired Dataset
-Following previous works [17; 20; 21], we use three types of indoor rooms in 3D-FRONT [72] and preprocess the dataset by filtering some problematic samples, resulting in 4041 bedrooms, 813 living rooms and 900 dining rooms.
+Following previous works [16; 19; 20], we use three types of indoor rooms in 3D-FRONT [71] and preprocess the dataset by filtering some problematic samples, resulting in 4041 bedrooms, 813 living rooms and 900 dining rooms.
 The number of objects $N_{i}$ in the valid bedrooms is between 3 and 12 with 21 object categories, i.e., $K_{c}=21$.
@@
   3. 3.
-    Caption Refinement: As generated captions may not always be accurate, we filter them with corresponding ground-truth categories using ChatGPT [28; 29], a large language model fine-tuned for instruction-based tasks.
+    Caption Refinement: As generated captions may not always be accurate, we filter them with corresponding ground-truth categories using ChatGPT [27; 28], a large language model fine-tuned for instruction-based tasks.
 This results in precise and expressive descriptions of each object in the scene.
@@
   4. 4.
-    Instruction Generation: The final instructions for scene synthesis are derived from $1\sim 2$ randomly selected “(subject, relation, object)” triplets obtained during the first extraction process.
+    Instruction Generation: The final instructions for scene synthesis are derived from 121$\sim$ 21 ∼ 2 randomly selected “(subject, relation, object)” triplets obtained during the first extraction process.
 Verbs and conjunctions within sentences are also randomly picked to maintain diversity and fluency.
@@
 5.1 for inpainting final posters, we obtained clean product images of significantly higher quality compared to the previous method.
-The comparisons between previous works [10; 12] are demonstrated in Figure 5.
+The comparisons between previous works [8; 10] are demonstrated in Figure 6.
 We provide an example in Figure 7 to showcase various elements contained in our curated poster dataset.
+A.3 Relation Definations
+Assume $X$ and $Y$ span the ground or poster plane, and $Z$ is the vertical axis for 3D scenes only.
+We use Center to represent the coordinates of a bounding box’s center.
+Height is the $Z$-axis size of a 3D bounding box.
+Relative orientation is computed as $\theta_{so}\coloneqq\text{atan2}(Y_{s}-Y_{o},X_{s}-X_{o})$, where $s$ and $o$ respectively refer to “subject” and “object” in a relationship.
+$d(s,o)$ is the ground distance between $s$ and $o$.
+$\text{Inside}(s,o)$ indicates whether the subject center is inside the ground bounding box of the object.
+We define 11 relationships in both 3D and 2D space as listed in Table IX and Table IX, respectively.
+A.4 Caption Refinement by ChatGPT
+The generated 3D object captions from BLIP are refined by ChatGPT (gpt-3.5-turbo).
+Our prompt to ChatGPT is provided in Table X.
+We set the hyperparameter temperature and top_p for text generation to 0.2 and 0.1 respectively, encouraging more deterministic and focused outputs.
+Appendix B Computational Costs
+Training takes about 2 days on one NVIDIA A40 GPU for the 3D bedroom and 2D poster datasets, and 3 days for the dining room and living room datasets, due to the increased complexity of semantic graphs in larger scenes.
+For inference, under the default configuration (diffusion steps $T=100+10$), generating a batch of 128 samples takes approximately 12 seconds for both 2D and 3D layout synthesis.
+Notably, our method can be significantly accelerated by reducing the number of diffusion steps.
+For example, setting $T=25+10$ reduces runtime to just 3 seconds, with minimal performance degradation.
+The impact of the diffusion step count is analyzed in Figure 4 and Table V.
+We believe that further acceleration is possible through advanced techniques such as flow matching and consistency distillation.
+Appendix C User Study
+For the evaluation of both 3D scene and 2D poster layout synthesis, a total of 360 questionnaires were gathered from 20 participants.
+This comprised 200 questionnaires for 3D scene layout synthesis and 160 for 2D poster layout synthesis.
+In each instance, participants were tasked with selecting the optimal synthesis result based on its fidelity to the given instructions.
+A method’s User Choice (UC) score is defined as the percentage of times its output was deemed superior by the participants, and the UC scores for all compared methods sum to 100% consequently.
+As shown in Figure 5, the empirical results demonstrate that InstructLayout achieved markedly higher preference rates over baseline methods, with user preference scores of 87.1% for 3D scene synthesis and 89.5% for 2D poster layout synthesis tasks, respectively.
+Figure 5: User study on layout synthesis tasks, encompassing (a) 3D scene generation and (b) 2D poster design.
+Appendix D Qualitative Results
+D.1 Instruction-driven 3D Scene Synthesis
+We present visualizations of instruction-driven synthesized 3D scenes in Figure 8, 9 and 10.
+Besides the quantitative evaluations shown in Table I, these qualitative visualizations also show the superiority of our method over previous state-of-the-art approaches in terms of adherence to instructions and generative quality.
+D.2 Zero-shot 3D Scene Applications
+We present visualizations illustrating various zero-shot instruction-driven applications, including stylization, re-arrangement, completion, and unconditional 3D scene synthesis in Figure 11, 12, 13 and 14 respectively.
+We find that the autoregressive model ATISS [16] tends to generate redundant objects, resulting in chaotic synthesized scenes.
+DiffuScene [20] encounters challenges in accurately modeling object semantic features, often yielding objects that lack coherence in terms of style or pairing, thereby diminishing the aesthetic appeal of the synthesized scenes.
+Moreover, both baseline models frequently struggle to follow the provided instructions during conditional generation.
+D.3 Instruction-driven 2D Poster Synthesis
+Comparisons of our synthesized 2D posters with other state-of-the-art methods are shown in Figure 15, 16, and 17, demonstrating the superiority of our method compared to previous state-of-the-art approaches in terms of better mimicking the behavior of human designers and following provided instructions.
+Appendix E Discussions
+Out-of-distribution Inputs.
+For out-of-distribution (OOD) inputs, the proposed system can still produce reasonable layouts, though they may not align closely with the input semantics.
+For instance, given instructions like “a black cat” or “a football playground”, the generated 3D indoor scenes tend to adopt black tones or football-themed decorations, while still maintaining a plausible bedroom layout, rather than generating a cat or an outdoor playground.
+Similarly, in 2D poster generation, when provided with natural images, the system interprets them as poster backgrounds and places plausible text bounding boxes accordingly.
+This generalization ability stems from the pretrained CLIP text and image encoders, which are adapted to extract semantically meaningful features from diverse inputs.
+Failure Cases.
+Although the proposed generative framework demonstrates strong performance and robustness to OOD inputs across both 3D indoor scenes and 2D posters, it still exhibits several failure cases.
+For example, the predicted 3D bounding boxes do not fully account for the complex relationships between meshes, leading to occasional object collisions in generated 3D scenes, which is also partly due to noisy annotations in the 3D-FRONT dataset.
+In the 2D poster generation task, the model sometimes produces redundant bounding boxes, especially when the background images already contain text.
+This can result in visually cluttered designs, as the generated boxes may unintentionally overlap with existing visual elements.
+Visualization results for these failure cases are provided in Figure 19.
+Limitation and Future Work.
+Another limitation of our work lies in its insufficient generalization across categories for both 3D indoor scene and 2D poster generation.
+The current approach relies on training a small, specialized generative model for each data type, whereas recent trends favor unified models trained on large-scale, diverse datasets.
+In light of the rapid development of large vision language models (VLMs), integrating a VLM into our instruction-driven pipeline holds strong potential for improving both generalization and controllability in future work [79; 80].
+Meanwhile, designing a unified framework that can generate either a 3D scene [81] or a 2D poster [82] in a single framework is also a promising direction, as it could simplify the pipeline design and accelerate inference.
 TABLE VII: Comparison of the types of elements contained in various poster datasets.
@@
 Figure 7: An example from our comprehensive 2D E-commerce poster dataset.
-A.3 Relation Definations
-Assume $X$ and $Y$ span the ground or poster plane, and $Z$ is the vertical axis for 3D scenes only.
-We use Center to represent the coordinates of a bounding box’s center.
-Height is the $Z$-axis size of a 3D bounding box.
-Relative orientation is computed as $\theta_{so}\coloneqq\text{atan2}(Y_{s}-Y_{o},X_{s}-X_{o})$, where $s$ and $o$ respectively refer to “subject” and “object” in a relationship.
-$d(s,o)$ is the ground distance between $s$ and $o$.
-$\text{Inside}(s,o)$ indicates whether the subject center is inside the ground bounding box of the object.
-We define 11 relationships in both 3D and 2D space as listed in Table IX and Table IX, respectively.
 TABLE VIII: Rules to determine the spatial relationships between objects in a 3D scene.
@@
 Closely below $-\frac{3\pi}{4}\leq\theta_{so}<-\frac{\pi}{4}$ and $d(s,o)\leq 80$
-Enclosing Inside($o$,$s$)
-Enclosed with Inside($s$,$o$)
+[…]
+[…]
 None $d(s,o)>300$
-A.4 Caption Refinement by ChatGPT
-The generated 3D object captions from BLIP are refined by ChatGPT (gpt-3.5-turbo).
-Our prompt to ChatGPT is provided in Table X.
-We set the hyperparameter temperature and top_p for text generation to 0.2 and 0.1 respectively, encouraging more deterministic and focused outputs.
 TABLE X: Prompt for ChatGPT to refine raw 3D object descriptions.
@@
 The old descriptions is ‘{BLIP caption}’, its category is ‘{ground-truth category}’, the new descriptions should be:
-Appendix B Qualitative Results
-B.1 Instruction-driven 3D Scene Synthesis
-We present visualizations of instruction-driven synthesized 3D scenes in Figure 7, 8 and 9.
-Besides the quantitative evaluations shown in Table I, these qualitative visualizations also show the superiority of our method over previous state-of-the-art approaches in terms of adherence to instructions and generative quality.
 (a) Instructions
@@
 Figure 10: Visualizations for instruction-drive synthesized 3D dining rooms by ATISS [16], DiffuScene [20] and our method.
-B.2 Zero-shot 3D Scene Applications
-We present visualizations illustrating various zero-shot instruction-driven applications, including stylization, re-arrangement, completion, and unconditional 3D scene synthesis in Figure 10, 11, 12 and 13 respectively.
-We find that the autoregressive model ATISS [17] tends to generate redundant objects, resulting in chaotic synthesized scenes.
-DiffuScene [21] encounters challenges in accurately modeling object semantic features, often yielding objects that lack coherence in terms of style or pairing, thereby diminishing the aesthetic appeal of the synthesized scenes.
-Moreover, both baseline models frequently struggle to follow the provided instructions during conditional generation.
 (a) Instructions
@@
 Figure 14: Visualizations for unconditional 3D scenes stylization by ATISS [16], DiffuScene [20] and our method.
-B.3 Instruction-driven 2D Poster Synthesis
-Comparisons of our synthesized 2D posters with other state-of-the-art methods are shown in Figure 14, 15, and 16, demonstrating the superiority of our method compared to previous state-of-the-art approaches in terms of better mimicking the behavior of human designers and following provided instructions.
 (a) Instruction & Product Information
@@
 Note that posters are rendered according to synthesized graphic and textual features.
+(a) Failure cases of 3D scene synthesis: object collisions resulting from limited modeling of inter-object relationships and noisy dataset annotations.
+(b) Failure cases of 2D poster synthesis: redundant bounding boxes and overlaps with existing background text due to insufficient content awareness.
+Figure 19: Failure cases of InstructLayout.